OpenStudy (anonymous):

What is the solution to the equation? 4x-26-3x=16-20 will give medals

pooja195 (pooja195):

Combine like terms :) 4x-3x=?

OpenStudy (anonymous):

7x

pooja195 (pooja195):

4x−26−3x=16−20 4x+−26+−3x=16+−20 (4x+−3x)+(−26)=(16+−20)(Combine Like Terms) x+−26=−4 x−26=−4 Add 26 to both sides. x−26+26=−4+26 x=?

OpenStudy (anonymous):

-30

pooja195 (pooja195):

26-4=?

OpenStudy (anonymous):

22

pooja195 (pooja195):

mhm :) \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

thank you @pooja195

pooja195 (pooja195):

:)